For hydrogen, an electron is in 2p state. Obtain the magnitude of orbital angular momentum and possible



z-components of orbital angular momentum 𝐿⃗ . What would be the possible values of total angular



momentums (𝐽 ) and associated magnetic moments 𝜇𝐽



⃗⃗⃗

1
Expert's answer
2022-01-12T08:32:08-0500

For P

State

l=1,s="\frac{1}{2}"

Calculate J value

L-s to L+s

Total anguler momentum

"J=\frac{1}{2},\frac{3}{2}"

Z- components of anguler momentum

ml= "\frac{-3}{2},\frac{-1}{2},\frac{1}{2},\frac{3}{2}"

"\\J=L-\frac{1}{2}"

We know that formula total angular moment




"\mu=g_{lp}(j-\frac{1}{2})+\frac{g_{sp}}{2}"

"g_{lp}=1,g_{sp}=5.585\mu_N"

"\mu=1\times(\frac{3}{2}-\frac{1}{2})+\frac{5.585}{2}"

total angular



momentum (𝐽 ) and associated magnetic moment


"\mu=1+2.7925=3.7925\mu_N"

"J=L-\frac{1}{2}"


"\mu=\frac{j}{j+1}(g_{lp}(j+\frac{3}{2})-\frac{g_{sp}}{2})"

"\mu=\frac{\frac{1}{2}}{\frac{1}{2}+1}(1\times(\frac{1}{2}+\frac{3}{2})-\frac{5.585}{2})"



"\mu=\frac{1}{3}(2-\frac{5.585}{2})=-0.264\mu_N"